05142008:

47 is the result when 12 and x are added. What is the value of x?

05142008:

thank u 🙂

jhonyy9:

47 = 12 +x x = ? how you get the value of x ? so to eliminate the 12 you need subtract the 12 from both sides 47-12 = 12-12 +x so x = ? hope helped understandably